. A motorcycle of mass 100 kilograms travels around a flat, circular track of radius  10 meters with a constant speed of 20 meters per second. What force is required  to keep the motorcycle moving in a circular path at this speed?



1
Expert's answer
2022-02-07T08:42:55-0500

The centripetal acceleration is given by

"a_{c}=\dfrac{V^{2}}{r}"

Wherein.

"V=20\;m/s" is the speed.

"r=10\;m" Is the acceleration.


Evaluating numerically

"a_{c}=\dfrac{V^{2}}{r}\\
a_{c}=\dfrac{(20\;m/s)^{2}}{10\;m}\\
a_{c}=40\;m/s^{2}"


The centripetal force is given by

"F_{C}=m\;a_{c}"

In where.

"a_{c}=40\;m/s^{2}" Is the centripetal acceleration.

"m=100\;kg" Is the mass.


Evaluating numerically.

"F_{C}=m\;a_{c}\\
F_{C}=100\;Kg\times 40\;m/s^{2}\\
F_{c}=4000\;N"


 The force required to keep the motorcycle moving in a circular path at this speed is

"F_{c}=4000\;N"